A rocket test sled accelerates at a constant rate of 15 m/s2 from rest while moving along a straight track. The rocket engine turns off instantly when the sled reaches a speed 30 m/s. The sled then moves freely (acceleration = 0) for a period of 5 s, after which the brakes are automatically applied giving a constant deceleration of -25 m/s2. What is the total time that the sled is moving?
Answer #1

Using 1st equation of motion,

V = u + a t

30 = 0 + 15 x t1

Time, t1 = 2 m/s^2

V = u + at

0 = 30 - 25 t2

Time, t2 = 1.2 sec

Total time, t = 1.2 + 2 + 5 = 8.2 sec

Comment in case any doubt please rate my answer....